# Extreme events in locally coupled bursting neurons

### v1

@@ II.3 Spiking to Bursting Transition
Figure 4: Time series (row 1), phase portrait (row 2), Interspike interval (ISI) distribution (row 3) and joint ISI histograms (row 4) of the collective observables of system (2) for the coupling strengths $k=-0.4$ (column 1), $k=-0.1$ (column 2), $k=0.1$ (column 3) and $k=0.4$ (column 4).
@@
Upon examining the dynamics of the system (2), we can understand that on changing the coupling strength $k$, dynamics of the collective observable $X_{m}$ make a transition from the spiking to bursting. To confirm this, we plot in Fig. 4, the time series (Row 1), phase portrait (Row 2), interspike interval (ISI) distribution (Row 3) and joint ISI histograms (Row 4). This analysis is carried out and portrayed for four different values of the coupling strength, say $k=-0.4$ (Column 1), $k=-0.1$ (Column 2), $k=0.1$ (Column 3) and $k=0.4$ (Column 4). These four values of $k$ are considered at $I=4.0$ from the two parameter probability plot (see red cross marks in Fig. 2(a)) Initially, for the coupling strength $k=-0.4$, from the time series (Fig. 4(a)) and phase portrait (Fig. 4(b)), we can understand that the system exhibits a four periodic spiking behaviour. Correspondingly, in the PDF plot (Fig. 4(c)), we can observe only 4 lines which confirm it is four-periodic. This is further confirmed by four dots in the joint ISI histogram (Fig. 4(d)). Next for $k=-0.1$, observing the time series (Fig. 4(e)) and phase portrait (Fig. 4(f)), we can infer that at this zone the system exhibits chaotic spiking behaviour. This coupling strength lies in the EE emerging zone. The PDF plot (Fig. 4(g)) resembles a Gaussian-type distribution which indicates the spiking is tonic [54]. This is also confirmed by the joint ISI histogram (Fig. 4(h)) where the joint ISI histogram forms an inverted L shape near the top right corner. Further, as we move down to the coupling strength $k=0.1$, the time series (Fig. 4(i)) shows combined chaotic spiking and bursting. Phase portrait (Fig. 4(j)) also confirms the mixture of these two dynamics. In particular, a small tail in the phase portrait ensures the presence of chaotic bursting as well. This is confirmed by ISI distribution (Fig. 4(k)) which looks similar to the bursting form even though there is a small Gaussian peak which shows that the dynamics have alternative chaotic bursting and spiking. In the joint ISI histogram (Fig. 4(l)), we can observe the same pattern as in the previous coupling strength. We can notice that in the EE zones, the patterns are similar. Finally, for $k=0.4$, we can observe purely bursting property (Fig. 4(m)) which is slightly chaotic and the phase portrait (Fig. 4(n)) confirms this as bursting. ISI PDF (Fig. 4(o)) and joint ISI histogram (Fig. 4(p)) also confirm that the dynamics are bursting type. In the ISI pdf plot, we can notice there are larger spike intervals between 0 and 20 and after a gap, we can notice spike intervals between 40 and 60. The interval between 0 and 20 belongs to the spike interval between spikes within bursts. Whereas the interval between 40 and 60 represents the interburst interval. Correspondingly, when we observe the joint ISI histogram (Fig. 4(p)), we can see the intervals spread all over the regions confirming the bursting property.
@@
In the above, we have demonstrated that system (2) undergoes a transition from spiking to bursting for increasing coupling strength $k$. Specifically, EE occur in two regions, namely regions of chaotic spiking (column 2 in Fig. 4) and regions of alternate spiking and bursting (column 3 in Fig. 4). Now we intend to investigate the mechanism by which system (2) gives birth to EE. Before proceeding further, it is important to note that due to the choice of the initial conditions, odd neurons with one set of IC’s get synchronized separately and form one cluster. While the even neurons with another set of IC’s get synchronized separately and form another cluster. At any point in time, the formation of two clusters is apparent. Due to the property of synchronous evolution of neurons inside a single cluster, we choose one neuron from each cluster and study the nature of time series from the two neurons alone. This will allow us to understand the dynamics between the clusters easily. So, we choose neuron $1$ from the odd cluster and neuron $2$ from the even cluster and name them henceforth as $x_{o}$ and $x_{e}$ respectively.
III Birth of Extreme Events: Mechanism
@@
To understand the mechanism, from the probability plot in Fig. 2(b), we consider three cases of the coupling strength, namely (i) $k=-0.4$, (ii) $k=-0.17$, and (iii) $k=0.4$ and recognize the mechanism by discerning the dynamics (at the micro level) between EE and non-EE regions. By doing so we will be able to know, what phenomena at the micro level is causing EE at the macro level.
@@
(i) $k=-0.4$: To understand in detail about the collective and individual dynamics, we plot in Fig. 5 the dynamics of the system (2) at $k=-0.4$. Specifically, Fig. 5(a) is the time series of the collective observable $X_{m}$, Fig. 5(b) represents the synchronisation error plot of odd and even neurons, Fig. 5(c) represents the time series of the individual neuron $x_{o}$ (blue solid line) overlapped with the time series of $x_{e}$ (orange dotted lines) and Fig. 5(d) is the corresponding spatio-temporal plot. Time series in Fig. 5(a), has no trajectories in the collective observable $X_{m}$ crossing the threshold $H_{s}=m+6\sigma=1.21$ line (dotted red horizontal line). If we carefully look at the time series plot of the individual clusters $x_{o}$ and $x_{e}$ (Fig. 5(c)), we can notice that cluster $x_{o}$ bursts exactly in antiphase with the cluster $x_{e}$. Here odd neurons burst when the even neurons are at rest and vice versa. This is apparent from Fig. 5(c). Further, both the clusters exhibit bursting with four spikes within each burst. This along with the antiphase evolution is the reason why the collective observable $X_{m}$ exhibits a four periodic spiking trajectory. Figure 5(b) gives us the overview of the dynamics in the ($X_{o}-X_{e}$) phase plane. We can see that no trajectories are in the transverse direction which is nothing but the synchronization manifold. This also confirms the antiphase synchronization occurring between them. Because of this, no trajectories of the collective observable $X_{m}$ cross the threshold. So, the probability of getting EE is zero. Next from the spatio-temporal plot in Fig. 5(d), we can spot the oscillations of each and every neuron. Here we can easily notice that odd neurons and even neurons are not synchronously evolving in time. At any given point in time, we can notice that the plot has alternate green dots and violet zones, that is a discontinuity in those green dots in the vertical direction. The green dots are the excitability of odd oscillators and the violet zones are the resting state of even oscillators or the reverse. This confirms the lack of synchronization (antiphase) between the odd and even neurons.
@@
(ii) $k=-0.17$: Next, we move on to the region between $k=-0.23$ and $k=-0.06$ where we notice the evolution of EE in the collective observable $X_{m}$. To observe the dynamics of the system in this region, we choose $k=-0.17$. Similar to the previous case, here also we plot similar four different plots in Fig. 6 and analyze the dynamics. Upon inferring Fig. 6(a), we can notice that the system exhibits chaotic spiking dynamics in the time series of $X_{m}$ and in the middle, rarely spikes are large enough to cross the threshold $H_{s}=m+6\sigma=1.23$. These rare occurrences of peaks crossing the $H_{s}$ (peak crossing the solid red-dashed line) are known as EE. Here a chaotic spiking scenario occurs which confirms that within this zone of the coupling strength, $X_{m}$ is chaotic. When we observe the dynamics of the two clusters ($x_{o}$ and $x_{e}$) in Fig. 6(c), we can notice that both the clusters exhibit a chaotic bursting property and the number of spikes within bursting is also chaotic, irregular and changes continuously from time to time. Importantly, they are out of phase with each other (blue solid line out of phase with orange dashed line in Fig. 6(c)). But when we look at the time series of the individual clusters at the time of EE in $X_{m}$, to our surprise, we find that clusters phase synchronize with each other (in Figs. 6(a) and 6(c)). Intermittently synchronization occurs between them, i.e., most of the time the clusters $x_{o}$ and $x_{e}$ are out of phase and only once in a while they phase synchronize with each other. Interestingly, whenever intermittent phase synchronization occurs between $x_{o}$ and $x_{e}$, an EE is found to occur in the $X_{m}$ observable. Even for a long range of the time series, rare emergence of EE is visible. From this, we can conclude that the single neuron exhibit chaotic bursting and collective observable exhibit chaotic spiking. Figure 6(b) shows us the projection of the antiphase synchronisation manifold of $x_{o}$ and $x_{e}$ neurons. From the phase plane, we can observe that the clusters are in antiphase state most of the time, but occasionally they are in synchronized state and EE emerge. In this phase plane, the trajectories in the transverse direction are very sparse confirming the rarity in the phase synchronization. To validate the phase synchronization among all the neurons, during the emergence of EE, we plot the spatio-temporal plot in Fig. 6(d). During all the time, odd and even clusters are not synchronized. But, at the time of the EE (peak crossing red dashed $H_{s}$ line in Fig. 6(a)), correspondingly in the spatio-temporal plot (6(d)), we can observe that all the neurons are phase synchronizing with each other, where violet regions and green dots correspond to the resting and bursting states respectively. At the times of EE, we can notice a strong green line showcasing the synchronization among the neurons.
@@
(iii) $k=0.4$: Next we choose another non-extreme region between $k=-0.5$ to $k=0.5$, say $k=0.4$ and try to understand the dynamics. So we plot in Fig. 7(a), the time series of the collective observable $X_{m}$. We can observe that the collective dynamics of neurons shift from chaotic spiking to bursting property with five within-burst spiking. No peaks cross the threshold and hence no EE. The collective exhibits bursting. The time series of the individual clusters in Fig. 7(c), shows the synchronization between $x_{o}$ and $x_{e}$ neurons. The amplitude may not match but they are in complete phase synchrony. No EE is evident in this region because as all the neurons get synchronized, the threshold value raises to a very high level. The synchronization of the odd and even neurons can also be observed from Fig. 7(b). Since at all times, the trajectories are well synchronised, it moves only in the transverse direction. Finally, Fig. 7(d) shows pretty much clearly how all the neurons get synchronized in most of the regions which clearly explains why we are not getting EEs in Fig. 7(a). Here we can spot a bright green lines throughout the time which confirms the synchronization.
Thus on the route to synchronization, there are regions of the coupling strength where intermittent synchronization occur among the clusters. During such an intermittent phase synchronization, EE emerge in the collective observable.
@@ IV Robustness of Cluster synchronization
Due to the choice of ICs, intermittent two cluster synchronizations are found to happen in the system (2) and as an effect, the system collectively is found to exhibit EE. In other words, we used two different initial conditions due to which system (2) evolves as a two-body problem i.e., odd neurons and even neurons got synchronised separately. Intermittent synchronization of these clusters leads to the emergence of EE in the chaotic region. Now let us discuss about the robustness of this cluster synchronization with respect to the increase in the number of different initial conditions. We have found that when the number of different initial conditions given to the system increases, correspondingly the number of clusters that the system splits into also increases. To prove this, we demonstrate below the phase synchronization of clusters using the spatio-temporal plots.
Figures 8(a), 8(b) and 8(c) show the dynamics of system (2) respectively for 4, 8 and 10 ICs. Subplots with index 1 $(a1,b1,c1)$ are the time series and subplots with index 2 $(a2,b2,c2)$ are spatio-temporal plots. First for the 4 ICs case, by observing Figs. 8(a1) and 8(a2) at the EE occurring part, three out of four clusters get involved in synchronization without the involvement. Next for 8 ICs, in Figs. 8(b1) and 8(b2) at the EE region, only five clusters get synchronized whereas three clusters are not synchronized. Finally for the 10 ICs case, in Figs. 8(c1) and 8(c2), at the EE point, we can observe that only seven clusters get synchronized while three clusters do not get synchronized. In addition, we have also checked for 100 different ICs from random generation. We find that in this case, if 25-30 oscillators come in a very small time zone, EE occur in the collective variable.
@@
As the number of clusters increases the occurrence of EE will become even more rarer. This is because intermittent synchronisation between the clusters will become rarer and rarer. This can be inferred from the IEI plots in Fig. 9 where the IEI distribution is fitted with generalized extreme value distribution (GEV). Initially for two ICs, we can notice the maximum time interval between the two EE will be 30,000 (see Fig. 3(b)). Next for 5 ICs, from Fig 9(a), we can observe that the maximum time interval is $16\times 10^{4}$. Figure 9(b) is the IEI for 10 different ICs where the maximum time interval is $70\times 10^{4}$. This shows that the interval between two EE increases as the number of different ICs increases. Similarly, for 100 ICs case, in the entire $10^{9}$ iterations, only 3 EE were found to occur. To explain this, we plot in Fig. 10, the number of EE peaks versus the number of ICs. In Fig. 10, $x$ and $y$ axes denote the number of ICs and number of EE respectively. Legend blue star represents the number of EE calculated numerically while the green dashed curve denotes the power law fit. The expression for the fit and the corresponding coefficient of determination $R^{2}$ values are given in the figure itself. From the plot, we can infer decrease in EE with an increase in ICs. That is ICs and EE are inversly proportional. This confirms the raise in the rarity of the EE with the number of number of ICs.
Figure 9: IEI distribution of system (2) at $k=-0.17$ for (a) 5 and (b) 10 different ICs.

### v2

@@ II.3 Spiking to Bursting Transition
Figure 6: Time series (row 1), phase portrait (row 2), Interspike interval (ISI) distribution (row 3) and joint ISI histograms (row 4) of the collective observable of the system (2) for the coupling strengths $k=-0.4$ (column 1), $k=-0.1$ (column 2), $k=0.1$ (column 3) and $k=0.4$ (column 4).
@@
Upon examining the dynamics of the system (2), we can understand that on changing the coupling strength $k$, dynamics of the collective observable $X_{m}$ make a transition from the spiking to bursting. To confirm this, we plot in Fig. 6, the time series (Row 1), phase portrait (Row 2), interspike interval (ISI) distribution (Row 3) and joint ISI histograms (Row 4). This analysis is carried out and portrayed for four different values of the coupling strength, in particular $k=-0.4$ (Column 1), $k=-0.1$ (Column 2), $k=0.1$ (Column 3) and $k=0.4$ (Column 4). These four values of $k$ are considered at $I=4.0$ from the two parameter probability plot (see red cross marks in Fig. 2(a)) Initially, for the coupling strength $k=-0.4$, from the time series (Fig. 6(a)) and phase portrait (Fig. 6(b)), we can understand that the system exhibits a four periodic spiking behaviour. Correspondingly, in the PDF plot (Fig. 6(c)), we can observe only 4 lines which confirm it is four-periodic. This is further confirmed by four dots in the joint ISI histogram (Fig. 6(d)). Next for $k=-0.1$, observing the time series (Fig. 6(e)) and phase portrait (Fig. 6(f)), we can infer that at this zone the system exhibits chaotic spiking behaviour. This coupling strength lies in the EE emerging zone. The PDF plot (Fig. 6(g)) resembles a Gaussian-type distribution which indicates the spiking is tonic [62]. This is also confirmed by the joint ISI histogram (Fig. 6(h)) where the joint ISI histogram forms an inverted L shape near the top right corner. Further, as we move down to the coupling strength $k=0.1$, the time series (Fig. 6(i)) shows combined chaotic spiking and bursting. Phase portrait (Fig. 6(j)) also confirms the mixture of these two dynamics. In particular, a small tail in the phase portrait ensures the presence of chaotic bursting as well. This is confirmed by ISI distribution (Fig. 6(k)) which looks similar to the bursting form even though there is a small Gaussian peak which shows that the dynamics have alternative chaotic bursting and spiking. In the joint ISI histogram (Fig. 6(l)), we can observe the same pattern as in the previous coupling strength. We can notice that in the EE zones, the patterns are similar. Finally, for $k=0.4$, we can observe purely bursting property (Fig. 6(m)) which is slightly chaotic and the phase portrait (Fig. 6(n)) confirms this as bursting. ISI PDF (Fig. 6(o)) and joint ISI histogram (Fig. 6(p)) also confirm that the dynamics is bursting type. In the ISI PDF plot (Fig. 6(o)), we can notice there are large probability for spikes having intervals between 0 and 20, followed by which there is a brief non-zero probability between 40 and 60 representing that there are no spikes present for that ISI. Then, there emerges a very tiny non-zero probability for the emergence of spikes intervaling between 40 and 60. Here, the non-zero probabilities for spike intervals ranging from 0 to 20 corresponds to the ISI within bursts. Whereas the very tiny non-zero probabilities for the spike intervals between 40 and 60 corresponds to the interburst ISI. Correspondingly, when we observe the joint ISI histogram (Fig. 6(p)), we can see the intervals spread all over the regions confirming the bursting property.
@@
In the above, we have demonstrated that the collective variable $X_{m}$ undergoes a transition from spiking to bursting for increasing coupling strength $k$. Specifically, EE occur in two regions, namely regions of chaotic spiking (column 2 in Fig. 6) and regions of alternate spiking and bursting (column 3 in Fig. 6). Now we intend to investigate the mechanism by which system (2) gives birth to EE. Before proceeding further, it is important to note that due to the choice of the initial conditions, odd neurons with one set of IC’s get synchronized separately and form one cluster and the even neurons with another set of IC’s get synchronized separately and form another cluster. At any point in time, the formation of two clusters is apparent. Due to the property of synchronous evolution of neurons inside a single cluster, we choose one neuron from each cluster and study the nature of time series from the two neurons alone. This will allow us to understand the dynamics between the clusters easily. So, we choose neuron $1$ from the odd cluster and neuron $2$ from the even cluster and name them henceforth as $x_{o}$ and $x_{e}$ respectively.
III Birth of Extreme Events: Local Mechanism
@@
To understand the mechanism, from the probability plot in Fig. 2(b), we consider three cases of the coupling strength, namely (i) $k=-0.4$, (ii) $k=-0.17$, and (iii) $k=0.4$ and recognize the mechanism by discerning the dynamics (microscopic) between EE and non-EE regions. By doing so we will be able to know, what phenomena at the micro level is causing EE at the macro level.
@@
(i) $k=-0.4$: To understand in detail about the collective and individual dynamics, we plot in Fig. 7 the dynamics of the system (2) at $k=-0.4$. Specifically, Fig. 7(a) is the time series of the collective observable $X_{m}$, Fig. 7(b) represents the synchronization error plot of odd and even neurons, Fig. 7(c) represents the time series of the individual neuron $x_{o}$ (blue solid line) overlapped with the time series of $x_{e}$ (orange dotted lines) and Fig. 7(d) is the corresponding spatio-temporal plot. Time series in Fig. 7(a), has no trajectories in the collective observable $X_{m}$ crossing the threshold $H_{s}=m+6\sigma=1.21$ line (dotted red horizontal line). If we carefully look at the time series plot of the individual clusters $x_{o}$ and $x_{e}$ (Fig. 7(c)), we can notice that cluster $x_{o}$ bursts exactly in antiphase with the cluster $x_{e}$. Here odd neurons burst when the even neurons are at rest and vice versa. This is apparent from Fig. 7(c). Further, both the clusters exhibit bursting with four spikes within each burst. This along with the antiphase evolution is the reason why the collective observable $X_{m}$ exhibits a four periodic spiking trajectory. Figure 7(b) gives us the overview of the dynamics in the ($x_{o}-x_{e}$) phase plane. We can see that no trajectories are in the transverse direction which is nothing but the synchronization manifold. This also confirms the antiphase synchronization occurring between them. Because of this, no trajectories of the collective observable $X_{m}$ cross the threshold. So, the probability of getting EE is zero. Next from the spatio-temporal plot in Fig. 7(d), we can spot the oscillations of each and every neuron. Here we can easily notice that odd neurons and even neurons are not synchronously evolving in time. At any given point in time, we can notice that the plot has alternate green dots and violet zones, that is a discontinuity in those green dots in the vertical direction. The green and violet regions are excitation and resting states, respectively. Alternate green and violet represents the spiking and resting states of alternate oscillators. This confirms the lack of synchronization (antiphase) between the odd and even neurons.
@@
(ii) $k=-0.17$: Next, we move on to the region between $k=-0.23$ and $k=-0.06$ where we notice the evolution of EE in the collective observable $X_{m}$. To observe the dynamics of the system in this region, we choose $k=-0.17$. Similar to the previous case, here also we plot similar four different plots in Fig. 8 and analyze the dynamics. Upon inferring Fig. 8(a), we can notice that the system exhibits chaotic spiking dynamics in the time series of $X_{m}$ and in the middle, large spikes rarely cross the threshold $H_{s}=m+6\sigma=1.23$. These rare occurrences of peaks crossing the $H_{s}$ (peak crossing the solid red-dashed line) correspond to EE. Here a chaotic spiking scenario occurs which confirms that within this zone of the coupling strength, $X_{m}$ is chaotic. When we observe the dynamics of the two clusters ($x_{o}$ and $x_{e}$) in Fig. 8(c), we can notice that both the clusters exhibit a chaotic bursting property and the number of spikes within bursts is irregular, and changes continuously from time to time. Importantly, they are out of phase with each other (blue solid line out of phase with orange dashed line in Fig. 8(c)). But when we look at the time series of the individual clusters at the time of EE in $X_{m}$, to our surprise, we find that clusters phase synchronize with each other (in Figs. 8(a) and 8(c)). Intermittent synchronization occurs between them, that is most of the time the clusters $x_{o}$ and $x_{e}$ are out of phase and only once in a while they phase synchronize with each other. Interestingly, whenever intermittent phase synchronization occurs between $x_{o}$ and $x_{e}$, an EE is found to occur in the $X_{m}$ observable. Even for a long range of the time series, rare emergence of EE is visible. From this, we can conclude that the single neuron exhibits chaotic bursting and collective observable exhibit chaotic spiking. The value of the largest Lyapunov exponent for the collective dynamics at $k=-0.17$ is found to be $1.0413$ confirming the chaotic nature of the dynamics. The corresponding Poincaré cross section at $k=-0.17$ is presented in Fig. 9. Data points only for $X_{m}>0$ are plotted. The sparse points on the right side corresponds to the extreme events.
Figure 9: Poincaré cross section of the collective dynamics of system 2 greater than $X_{m}>0$ at $k=0.17$. This confims the chaotic nature of the system. The sparse points at the right end of the plot represents the rare extreme events.
Figure 8(b) shows us the projection of the antiphase synchronization manifold of $x_{o}$ and $x_{e}$ neurons. From the phase plane, we can observe that the clusters are in antiphase state most of the time, but occasionally they are in synchronized state and EE emerge. In this phase plane, the trajectories in the transverse direction are very sparse confirming the rarity in the phase synchronization. To validate the phase synchronization among all the neurons, during the emergence of EE, we plot the spatio-temporal plot in Fig. 8(d). During all the time, odd and even clusters are not synchronized. But, at the time of the EE (peak crossing red dashed $H_{s}$ line in Fig. 8(a)), correspondingly in the spatio-temporal plot (Fig. 8(d)), we can observe that all the neurons are phase synchronizing with each other, where violet regions and green dots correspond to the resting and bursting states respectively. At the times of EE, we can notice a strong green line showcasing the synchronization among the neurons. It is also important to note that in the inhibitory coupling, the cluster synchronization is a phase synchronization, whereas when the coupling is excitatory, the cluster synchronization is complete synchronization.
@@
(iii) $k=0.4$: Next we choose another non-extreme region between $k=-0.5$ to $k=0.5$, namely, $k=0.4$ and try to understand the dynamics. So we plot in Fig. 10(a), the time series of the collective observable $X_{m}$. We can observe that the collective dynamics of neurons shift from chaotic spiking to bursting property with five within-burst spiking. No peaks cross the threshold and hence no EE. The collective observable $X_{m}$ exhibits bursting. The time series of the individual clusters in Fig. 10(c), shows the synchronization between $x_{o}$ and $x_{e}$ neurons. The amplitude may not match but they are in complete phase synchrony. No EE is evident in this region because as all the neurons get synchronized, the threshold value raises to a very high level. The synchronization of the odd and even neurons can also be observed from Fig. 10(b). Since at all times, the trajectories are well synchronized, it moves only in the transverse direction. Finally, Fig. 10(d) shows pretty much clearly how all the neurons get synchronized in most of the regions which clearly explains why we are not getting EEs in Fig. 10(a). Here we can spot a bright green lines throughout the time which confirms the synchronization.
Analysis done so far, reveals that for a unidirectional ring of HR neurons with clustered ICs, varying the coupling strength transits the oscillators present in the system from anti-phase to in-phase synchronization. This transition happens gradually with the occurrence of occasional in-phase synchronization in time. As the coupling strength is further increased, occasional in-phase synchronization becomes frequent and for higher values of the coupling strength, burst synchronize occurs between them. Thus on the route to synchronization, there are regions of the coupling strength where intermittent synchronization occur among the clusters. During such an intermittent phase synchronization, EE emerge in the collective observable. This intermittent phase synchronization between two cluster of oscillators (intermittent cluster synchronization) is the potential reason behind the emergence of EE in the considered system. It is well-known that systems with symmetries will undergo a transition from antiphase to in-phase synchronized state upon varying the coupling strength [52, 53, 54, 55, 56]. But in the present scenario, independent of the number of clusters of ICs, intermittent phase synchronization between the clusters occur leading to EE. Two things are important to be noted here. (i) All these results were obtained for $10^{9}$ iterations after leaving out $3\times 10^{5}$ iterations as transients. (ii) This phenomena is true even when the ICs was randomly picked from the phase portrait basin. We checked the dynamics for 100 realizations and for all, qualitatively the dynamics was found to be the same. These two properties confirm that the intermittent cluster synchronization phenomenon that take place in the system is a well settled structurally stable behavior and not a transient. Further, this phenomenon is found to be robust with respect to the network size. Further, it is to be noted that similar phenomenon is not observed for non-local or global network topology. In these cases, chaotic dynamics existed in the mean field observable, but no EE are observed.
@@ IV Robustness of Cluster synchronization
Due to the choice of ICs, intermittent two cluster synchronizations are found to happen in the system (2) and as an effect, the system collectively is found to exhibit EE. In other words, we used two clusters of initial conditions due to which system (2) evolves as a two-body problem that is odd neurons and even neurons got synchronized separately. Intermittent synchronization of these clusters leads to the emergence of EE in the chaotic region. Now let us discuss about the robustness of this cluster synchronization with respect to the increase in the number of clusters of initial conditions. We have found that when the number of clusters of initial conditions given to the system increases, correspondingly the number of clusters that the system splits into also increases. To prove this, we demonstrate below the phase synchronization of clusters using the spatio-temporal plots.
Figures 11(a), 11(b) and 11(c) show the dynamics of system (2) respectively for 4, 8 and 10 ICs. Subplots with index 1 $(a1,b1,c1)$ are the time series and subplots with index 2 $(a2,b2,c2)$ are spatio-temporal plots. First for the 4 ICs case, by observing Figs. 11(a1) and 11(a2) at the EE occurring part, three out of four clusters get involved in synchronization without the involvement. Next for 8 ICs, in Figs. 11(b1) and 11(b2) at the EE region, only five clusters get synchronized whereas three clusters are not synchronized. Finally for the 10 ICs case, in Figs. 11(c1) and 11(c2), at the EE point, we can observe that only seven clusters get synchronized while three clusters do not get synchronized. In addition, we have also checked for 100 different ICs from random generation. We find that in this case, if 25-30 oscillators come in a very small time zone, EE occur in the collective variable.
@@
As the number of clusters increases the occurrence of EE will become even more rarer. This is because intermittent synchronization between the clusters will become rarer and rarer. This can be inferred from the IEI plots in Fig. 12 where the IEI distribution is fitted with generalized extreme value distribution (GEV). Initially for two ICs, we can notice the maximum time interval between the two EE will be 30,000 (see Fig. 5(b)). Next for 5 ICs, from Fig 12(a), we can observe that the maximum time interval is $16\times 10^{4}$. Figure 12(b) is the IEI for 10 different ICs where the maximum time interval is $70\times 10^{4}$. This shows that the interval between two EE increases as the number of different ICs increases. Similarly, for 100 ICs case, in the entire $10^{9}$ iterations, only 3 EE were found to occur. To explain this, we plot in Fig. 13, the number of EE peaks versus the number of ICs. In Fig. 13, $x$ and $y$ axes denote the number of ICs and number of EE respectively. Legend blue star represents the number of EE calculated numerically while the green dashed curve denotes the power law fit. The expression for the fit and the corresponding coefficient of determination $R^{2}$ values are given in the figure itself. From the plot, we can infer decrease in EE with an increase in ICs. That is ICs and EE are inversely proportional. This confirms the raise in the rarity of the EE with the number of number of ICs.
Figure 12: IEI distribution of system (2) at $k=-0.17$ for (a) 5 and (b) 10 different ICs. Blue line is the histogram of the simulated data whereas red line is the fitted GEV distribution.